OpenStudy (anonymous):

medal and fan. Find a number whose product with 8 is the same as its sum with 49. Select one: a. 8 b. 7 c. 6 d. 9

OpenStudy (anonymous):

whats a product anyway?

OpenStudy (johnweldon1993):

Hmm let the number = x Product = Multiplication :)

OpenStudy (johnweldon1993):

So x * 8 = x + 49 Now can you solve for 'x' ?

OpenStudy (anonymous):

is it C ?

OpenStudy (johnweldon1993):

Not quite... 8x = x + 49 -x -x 7x = 49 x = ?

OpenStudy (anonymous):

Duhh. 7

OpenStudy (johnweldon1993):

Haha perfect :)

OpenStudy (anonymous):

gosh. i really screw myself up lol.

OpenStudy (johnweldon1993):

Just little mistakes, you get them right in the end :)

OpenStudy (anonymous):

yeahh.(; i supose

OpenStudy (johnweldon1993):

lol always ;D